hello can you please help me with these 4 questions please

If the Fed sells $50 billion of securities to banks and RRR = 0.1, money supply will A. Increase by $500 billion B. Decrease by $500 billion C. Increase by $50 billion D. Decrease by $50 billion 10.00000 points

QUESTION 24 Let P = 200, Y = 10 and M = 400. Velocity of money is A. V = 2 B. V = 3 C. V = 4 D. V = 5

QUESTION 25 When the economy grows at the rate of 3% a year, a 7% growth in money supply would result in a 10% inflation 10% deflation 4% inflation 4% deflation points

QUESTION 26 Let the amount of currency and coins = $0.5 trillion, demand deposits = $0.3, travelers checks = $0.2 trillion, saving deposits = $0.5 trillion, small denomination time deposits = $6 trillion and mutual fund deposits = $2 trillion. Narrow measure of money supply is M1 = $0.5 trillion M1 = $0.8 trillion M1 = $1 trillion M1 = $1.5 trillion

QUESTION 27 Let M1 = $1 trillion, saving deposits = $0.5 trillion, small denomination time deposits = $6 trillion and mutual fund deposits = $2 trillion. Broad measure of money supply is M2 = $7.5 trillion M2 = $8.5 trillion M2 = $9.5 trillion M2 = $10.5 trillion
Answer #1

23.

B

Selling of securities, will decrease the money supply.

Decrease in money supply = 50/.1 = $500 billion

24.

D

MV = PY

400*V=200*10

V=5

25.

A

26.

M1= $1trillion

M1=.5+.3+.2 = $1 trillion

27.

M2=$9.5 trillion

M2=1+.5+6+2=$9.5 trillions